### v3

@@ 3.1.1 Connection for AI
For emerging services like AI, the current network’s support in communication is yet to be clarified, necessitating enhanced design to ensure that future 6G networks can better meet the new demands of AI services from a connectivity perspective. Considering the control plane and user plane separation principle [121], the types of AI service-related data carried by communication connections can be naturally categorized into two major groups: AI signaling and AI data. Consequently, connection for AI embodies providing connectivity support for AI signaling and data.
@@
AI Signaling: 1) The AI Signaling is used for transmitting AI-related control messages, e.g., AI service request/response messages; request/response messages related to computing power required for AI analysis, etc. 2) For the transmission of AI analysis information, including intermediate analysis information (when multiple entities perform joint analysis).
@@
AI Data: 1) The AI data is used to transfer AI input data, generally referred to as input vectors for invoking an AI model or represented as inputs in an AI algorithm. 2) It is also used for transmitting AI outputs, generally called the output of the invoked AI model/algorithm. 3) It can be used for transmitting AI models, e.g., transmitting a neural network model usually means transmitting its parameter weights (which can be represented as vectors) and its network structure (which can be represented as scalars). 4) It can be further used for transmitting AI training data, which is usually huge and dedicated to the training phase of intelligent model.
@@
For AI business scenarios and related connectivity performance requirements, 3GPP TS 22.261 (based on the conclusions of TR 22.874) also has some KPIs requirement level analysis and results for 5G systems, mainly on the impact of conventional connectivity performance metrics, such as throughput, latency, reliability, etc., in other words, the current 3GPP standard only analyzes the impact of the 5G system (mainly the CN portion) in terms of how it supports different AI services/applications from the functional and performance perspectives. From an overall end-to-end process analysis perspective, AI signaling and AI data involved in AI services/applications necessitate transmission via the air interface. However, in the current standards and architectures, no relevant modules guide the transmission process or standardize the transmission content. As a result, AI services, inherently demand that the network provide better connection support for AI based on connection.
@@ 3.1.2 Data for AI
Future mobile communication networks will continuously generate massive amounts of data with diverse types, and data is precisely one of the core production factors for AI, providing a solid data foundation for 6G’s endogenous intelligent AI services [130]. In the field of communications, several publicly available datasets have been widely utilized, such as DeepMIMO [131] and RadioML [132]. The DeepMIMO dataset, based on ray-tracing simulations of real wireless communication environments, provides MIMO channel data and is extensively applied in scenarios like channel modeling, beamforming, and user location in 6G networks. Meanwhile, the RadioML dataset encompasses radio signal data with multiple modulation schemes and is mainly used for tasks including signal modulation recognition, interference detection, and spectrum management. Additionally, the “AI/ML in 5G Challenge” dataset [133] released by ITU covers a variety of communication scenarios in 5G networks, offering crucial support for research on network optimization, traffic prediction, and resource allocation. These datasets not only lay a solid foundation for AI research in the communications field but also provide a unified benchmark for validating and comparing algorithm performances.
Data for AI aims to efficiently support end-to-end data collection, transmission, storage, and sharing, addressing how to facilitate, expedite, and securely provide data to AI functions within or external to the 6G network. This endeavor strives to reduce the difficulty in data acquisition, enhance data service efficiency, and facilitate the realization of 6G AI while adhering to privacy and security laws and regulations. Depending on the potential functional scope, the data support provided by data for AI should encompass five components:
Data Collection/Distribution: This component provides a basic publish-subscribe mechanism for data producers and consumers, enhancing data collection/distribution efficiency and supplying samples for large-scale training required by AI models to achieve optimal results.
Data Security and Privacy: This component leverages security and privacy protection technologies to provide high-quality, trusted data services tailored to user and network needs, ensuring user and network privacy protection and data security, immutability, and traceability. By safeguarding data privacy, it also protects the security of AI models during training and data transmission.
Data Analysis: This component utilizes AI models, algorithms, and knowledge to provide statistical information, predictive insights, network anomaly analysis, and optimization suggestions, enhancing the data consumption experience for internal and external network functions and related intelligent applications.
Data Preprocessing: This component applies generic tools for format conversion, noise reduction, feature extraction, and other preprocessing tasks to collected data, fulfilling AI model training requirements for input data and facilitating the management of multi-dimensional complex data.
Data Storage: This component stores and retrieves collected data, providing storage support for data services related to data security and privacy, data analysis, or data preprocessing, and further establishing a data foundation to support AI services.
@@
Through standardized data service functions, a complete data service process can be established within the 6G network, providing a robust foundation for AI at the data level. Meanwhile, in view of the fact that researchers often need to construct dedicated datasets for specific application scenarios, Data for AI provides valuable guidance. Taking the construction of datasets in the field of communications as an example, we offer the following practical experiences. Firstly, data sources can be selected from collected data in real communication networks or generated simulation data using high-fidelity simulation tools (such as NS-3 or OMNeT++). Real data can reflect actual scenarios but may involve privacy and security issues; whereas simulation data offers higher flexibility and can be customized to specific scenarios according to particular requirements. Secondly, data preprocessing and labeling are essential steps. Communication data usually contains a large amount of noise, thus requiring preprocessing operations such as cleaning and normalization, and accurate labeling based on research objectives (such as channel state prediction or interference detection). Moreover, to enhance the generalization ability of the model, diverse communication scenarios and multiple network conditions should be covered as much as possible when constructing the dataset. Meanwhile, when using real user data, strict compliance with data privacy protection regulations is mandatory, and techniques such as anonymization or differential privacy should be adopted to ensure data security. Finally, we encourage researchers to make datasets publicly available under the premise of protecting privacy, in order to promote collaborative research and algorithm comparison within the field, and it is recommended to use a unified storage format and standard to facilitate data sharing and the reproducibility of research results.
@@ 3.1.3 Computation for AI
Computation is another core production factors of AI. In future network architectures, computational power will be ubiquitous across various infrastructures, including central clouds, edge clouds, network devices, and even end-user devices, providing a solid computational foundation for the diversification of AI services across all industries in 6G [134]. This computational power, along with the AI algorithms or functionalities attached, serves the network or devices to improve performance and optimize network operations but is also potentially exploited through unified interfaces to serve upper-layer applications. Consequently, computation for AI focuses on how the widely available computational power in the network can support AI services. This computational power can be broadly categorized into three types: dedicated computing power for network elements, distributed external computing power, and endogenous computing power within the network.
@@
Network Element Dedicated Computing Power (for AI): This type of computing power typically consists of computing and storage units composed of specialized processors and programmable devices. It serves network elements in mobile communication networks (BS or CNs). It is primarily used for customized AI application services (i.e., AI4NET) to enhance communication performance or optimize network operations, like CSI feedback, channel estimation, and beamforming. Due to resource restrictions, it cannot support AI services and applications requiring large-scale computations and training. It is challenging to implement for third-party applications.
@@
Distributed External Computing Power (for AI): This type of computing power typically exists in the form of distributed edge computing/MEC, primarily utilizing general-purpose central processing units (CPU) and potentially including high-performance graphics processing units (GPU) and programmable acceleration cards. External computing power shifts computations from centralized data centers to the edge of the access network, enabling network optimization and supporting industry applications with high computational demands and stringent latency requirements, such as video acceleration and augmented reality (AR)/VR scenarios. Due to its external nature, it requires management through unified network functions, introducing some latency that may impact extremely latency-sensitive AI services and applications. Moreover, this external overlay approach may not optimize resource utilization, hindering efficient construction, deployment, and usage of AI services.
@@
Distributed Network-Native Computing Power (for AI): In the vision of future networks, each network element will possess control, forwarding, and computing capabilities, with numerous computing nodes deployed throughout the network. This computational power, known as endogenous computing power, promotes the development and deployment of endogenous AI, enabling large-scale intelligent distributed collaborative services. It compensates for the shortcomings of external computing power, promptly responding to changes in mobility and networks while fostering the emergence and development of future AI applications, such as immersive cloud XR, holographic communication, sensory interconnection, communication-sensing integration, and DT.
Through the construction of multi-category computational power, 6G networks will provide abundant computing services, supporting diversified AI services and accelerating the training process of AI models.
@@ 3.1.4 Security for AI
Some security vulnerabilities remained after the design of 2G-5G systems, and conventional security measures often failed to ensure safety. In recent years, trusted endogenous security has emerged as a new security approach for 6G, characterized by four main features: collaboration, intelligent proactive defense, trustworthiness, and privacy protection. The concept of trusted endogenous security also applies to ensuring the security of the AI plane.
Ubiquitous Collaboration: Collaboration has become a technological approach to enhancing communication capabilities. Intelligent ubiquitous collaboration will enable NET4AI to adapt to various complex and dynamic environments and scenarios while improving the robustness of the AI plane to identify potential malicious attacks or abnormal behaviors accurately. Future security defenses will shift from isolated to highly efficient collaboration, achieving individual collaborative defenses between entities or intelligent agents, inter-layer defenses between protocol layers or architecture layers, and inter-domain defenses between network space domains.
Intelligent Active Defense: The rapid development of 6G networks and AI technology presents opportunities for intelligent security defenses. The intelligent, proactive defense will enable NET4AI to continuously learn and adapt while automatically adjusting defense strategies based on changes in the network environment and AI security threats to counteract attacks promptly. This transition represents a shift from passive protection to proactive sensing and defense [135].
Trustworthiness: 6G will further evolve towards trustworthiness capabilities, achieving a comprehensive endogenous security system encompassing Security (i.e., confidentiality, integrity, availability), reliability, resilience, and safety [136]. This can provide a trusted ecosystem for NET4AI. Furthermore, flexible and reliable access control [137] and effective identity authentication[138] can, to some extent, prevent unauthorized access from impacting the endogenous trusted environment of NET4AI.
Privacy Protection: The deep integration of AI, big data, and 6G networks will exacerbate the challenges of data privacy protection. NET4AI must handle massive business and personal data while preventing privacy leakage. Emerging distributed ML, while achieving distributed ubiquitous intelligence and partially avoiding privacy leakage caused by AI model training, still has security risks [139]. Therefore, NET4AI will integrate distributed ML with existing privacy protection methods at multiple points to enhance data privacy protection and build an efficient and secure data ecosystem.
@@ 3.1.5 Orchestration for AI
Orchestration for AI of the intelligence plane signifies efficient support for deploying, operating, and optimizing AI services within the network through orchestration. It automatically configures and orchestrates network resources based on the specific requirements of AI services, ensuring that AI applications obtain the necessary computational power, storage, network bandwidth, and other resources. Additionally, this orchestration capability enables real-time monitoring of AI service performance, dynamically adjusting resource allocation to accommodate fluctuations in service loads, thereby guaranteeing the stability and efficiency of AI services [140]. This seamless integration and support facilitates smoother AI service operation within the network, providing robust network guarantees for various intelligent applications. Specifically, this manifests in the following three aspects:
@@
Automated Deployment and Resource Allocation: The orchestration capability swiftly responds to AI services’ network resource demands through automation. Upon AI service deployment, the orchestrator intelligently analyzes the current network resource utilization, automatically configuring network devices, establishing links, and deploying services. This automated process enhances deployment efficiency, mitigates human errors, and ensures rapid AI service launch and stable operation.
@@
Dynamic Optimization and Elastic Scaling: As AI services operate, their resource requirements may fluctuate. The orchestration capability features dynamic optimization and continuous monitoring of AI service performance, including CPU usage, memory consumption, network bandwidth, and other key metrics. Upon detecting resource shortages or surpluses, the orchestration system automatically adjusts resource allocation, such as adding compute nodes, expanding storage capacity, or optimizing network bandwidth, to meet AI services’ dynamic demands. Furthermore, the orchestration system supports elastic scaling, automatically resizing resources based on AI service loads, ensuring service stability and efficiency.
@@
Fault Self-Healing and Intelligent O&M: By integrating monitoring, alerting, and troubleshooting functions, the orchestrator promptly identifies and locates faults within the network, automatically triggering recovery mechanisms. This fault self-healing capability minimizes service disruptions for AI services, enhancing service availability and reliability. Additionally, the orchestration facilitates intelligent network O&M, empowering network administrators to manage network resources and AI services more efficiently, reducing O&M costs, and improving efficiency.

### v4

@@ 3.1.1 Connection for AI
In 6G, the control plane and user plane will be further enhanced, aiming to deepen the service-oriented and open nature of network functions and provide support for emerging AI services. On the one hand, 6G needs to further enhance and optimize the service management and interface of the existing control plane functions to reduce the complexity of network functions and signaling interaction processes[134]. Additionally, 6G will extend the conventional control plane by integrating space-terrestrial ubiquitous access control management, enabling new functionalities and service-oriented offerings for session management and policy control across communication, computing, data, and other essential elements. On the other hand, the user plane will be enhanced by augmenting or introducing new, more generalized, and flexible user plane forwarding protocols that can perform agile service forwarding and processing based on communication, computing, and other elements. Leveraging microservice technologies and next-generation hardware support, e.g., high-performance programmable chips, cloud-native technologies, and the P4 (programming protocol-independent packet processors) advanced language, user plane functions (UPF) will be restructured into service-oriented modules[118, 135]. This flexible invocation and efficient collaboration among functional modules will break the constraints of conventional layered protocols, enhancing data forwarding efficiency and bandwidth utilization.
@@
Considering the control plane and user plane separation principle [126], the types of AI service-related data carried by communication connections can be naturally categorized into two major groups: AI signaling and AI data. Connection for AI embodies providing connectivity support for AI signaling and data.
@@
AI Signaling. AI Signaling is used for transmitting AI-related control messages, e.g., AI service request/response messages; request/response messages related to computing power required for AI analysis, etc. It can also include auxiliary information for collaborative AI analysis among multiple entities.
@@
AI Data. AI data is used to transfer AI input data, generally referred to as input vectors for invoking an AI model or represented as inputs in an AI algorithm. It is also used for transmitting AI outputs, generally called the output of the invoked AI model/algorithm. It can be used for transmitting AI models, e.g., transmitting a neural network model usually means transmitting its parameter weights (which can be represented as vectors) and its network structure (which can be represented as scalars). It can be further used for transmitting AI training data. The data can often be usually huge and dedicated to the training phase of intelligent models.
@@
For AI business scenarios and related connectivity performance requirements, 3GPP TS 22.261[136] (based on the conclusions of TR 22.874[137]) also has some KPIs requirement level analysis and results for 5G systems, mainly on the impact of conventional connectivity performance metrics, such as throughput, latency, reliability, etc.; in other words, the current 3GPP standard only analyzes the impact of the 5G system (mainly the CN portion) in terms of how it supports different AI services/applications from the functional and performance perspectives. From an overall end-to-end process analysis perspective, AI signaling and AI data involved in AI services/applications necessitate transmission via the air interface. However, in the current standards and architectures, no relevant modules guide the transmission process or standardize the transmission content.
@@ 3.1.2 Data for AI
6G networks necessitate a novel data capability that differs from the conventional user plane. This capability must adhere to data regulation and supervision requirements, while enhancing data analysis and processing efficiency, addressing data management challenges, and supporting AI services. This data capability supports the realization of information collection, cleaning, and storage, and enables the complete lifecycle management of data service processes while ensuring data privacy and security. By leveraging a hybrid data processing and storage model that combines centralized and distributed approaches, the 6G network can flexibly address diverse scenario requirements: 1) For sensitive or high-privacy-level data, local processing and storage can be performed at the RAN/CN and terminal sides to ensure no data leakage; 2) Global and comprehensively analyzed data, on the other hand, is processed and stored centrally in the data center. Through tight cooperation between the central data center and edge nodes, 6G can efficiently manage massive amounts of data originating from various technological and business domains while safeguarding privacy.
@@
Data capability provides a solid data foundation for 6G’s endogenous intelligent AI services [138]. Several publicly available datasets have been widely utilized, such as DeepMIMO [139] and RadioML [140]. The DeepMIMO dataset, based on ray-tracing simulations of real wireless communication environments, provides MIMO channel data and is extensively applied in scenarios like channel modeling, beamforming, and user location in 6G networks. By contrast, the RadioML dataset encompasses radio signal data with multiple modulation schemes and is mainly used for signal modulation recognition, interference detection, and spectrum management. Moreover, the “AI/ML in 5G Challenge” dataset [141] released by ITU covers a variety of communication scenarios in 5G networks, offering crucial support for research on network optimization, traffic prediction, and resource allocation. These datasets not only lay a solid foundation for AI research in the communications field but also provide unified benchmarks for validating and comparing algorithm performances.
Data for AI aims to efficiently support end-to-end data collection, transmission, storage, and sharing, addressing how to facilitate, expedite, and securely provide data to AI functions within or external to the 6G network. Depending on the potential functional scope, the data support provided by data for AI should encompass five components: 1) Data Collection/Distribution: This component provides a basic publish-subscribe mechanism for data producers and consumers, enhancing data collection/distribution efficiency and supplying samples for large-scale training required by AI models to achieve optimal results. 2) Data Security and Privacy: This component leverages security and privacy protection technologies to provide high-quality, trusted data services tailored to user and network needs, ensuring user and network privacy protection and data security, immutability, and traceability. 3) Data Analysis: This component utilizes AI models, algorithms, and knowledge to provide statistical information, predictive insights, network anomaly analysis, and optimization suggestions, enhancing the data consumption experience for internal and external network functions and related intelligent applications. 4) Data Preprocessing: This component applies generic tools for format conversion, noise reduction, feature extraction, and other preprocessing tasks to collected data, fulfilling AI model training requirements for input data and facilitating the management of multi-dimensional complex data. 5) Data Storage: This component stores and retrieves collected data, providing storage support for data services related to data security and privacy, data analysis, or data preprocessing, and further establishing a data foundation to support AI services.
@@
Through standardized data service functions, a comprehensive data service process can be established within the 6G network. As researchers often need to construct dedicated datasets for specific application scenarios, Data for AI can provide valuable guidance. Take the construction of communication datasets as an example. Firstly, data sources can be selected from collected data in real-world communication networks or generated simulation data using high-fidelity simulation tools (such as NS-3 or OMNeT++). Real-world data can reflect actual scenarios but may involve privacy and security issues; whereas simulation data offer higher flexibility and can be customized to specific scenarios according to particular requirements. Secondly, data preprocessing and labeling are two essential steps. Communication data usually contain a large amount of noise, thus requiring preprocessing operations, such as cleaning and normalization, and accurate labeling based on research objectives (e.g., channel state prediction or interference detection). When using real-world user data, strict compliance with data privacy protection regulations is mandatory. Techniques such as anonymization or differential privacy should be adopted. Last but not least, researchers are encouraged to make datasets publicly available under the premise of protected privacy, to promote collaborative research and algorithm comparison within the field. It is recommended to use a unified storage format and standard to facilitate data sharing and reproducibility of research results.
@@ 3.1.3 Computation for AI
The evolution of 6G will bring significant advancements in computation, including computational sensing, control, and execution. Computational sensing will gather information about the ubiquitous computing resources distributed across the spectrum, from central clouds to terminal devices. Computational control will cater to computing service demands by providing intelligent strategies for pervasive computing power control and resource scheduling. Computational execution, guided by these strategies, will deliver services such as session management, traffic control, and policy execution.
@@
Computational capabilities must synergize with communication connectivity to optimize the efficiency and energy consumption of emerging services like AI. This integration necessitates the development of a unified environment that ensures continuous operation of computing services across terminals, edge nodes, networks, and clouds. Such an environment will enable the dynamic selection of optimal computing nodes for computational tasks, considering factors like latency, bandwidth, computing power, and energy efficiency. This unified framework will allow computational tasks to transition among terminals, networks, edge nodes, and clouds[142]. This computational power, along with the AI algorithms or functionalities, serves the network or devices to improve network operations but is also potentially exploited through unified interfaces to serve upper-layer applications. Computational power can be categorized into three types: dedicated computing power for network elements, distributed external computing power, and endogenous computing power within the network.
@@
(1) Network Element Dedicated Computing Power for AI. This type of computing power typically consists of computing and storage units composed of specialized processors and programmable devices. It serves network elements in mobile communication networks (BS or CN). It is primarily used for customized AI application services (i.e., AI4NET) to enhance communication performance or optimize network operations, like CSI feedback, channel estimation, and beamforming. Due to resource restrictions, it cannot support AI services and applications requiring large-scale computations and training. It is challenging to support third-party applications.
@@
(2) Distributed External Computing Power for AI. This type of computing power typically exists in the form of distributed edge computing/MEC, primarily utilizing general-purpose central processing units (CPU), high-performance graphics processing units (GPU), and programmable acceleration cards. External computing power shifts computations from centralized data centers to the edge of the access network, enabling network optimization and supporting industry applications with high computational demands and stringent latency requirements, such as video acceleration and augmented reality (AR)/VR scenarios. Due to its external nature, it requires management through unified network functions, introducing some latency that may impact extremely latency-sensitive AI services and applications. Moreover, this external overlay approach may not optimize resource utilization, hindering efficient construction, deployment, and usage of AI services.
@@
(3) Distributed Network-Native Computing Power for AI. In the vision of future networks, each network element will possess control, forwarding, and computing capabilities, with numerous computing nodes deployed throughout the network. This computational power, known as endogenous computing power, promotes the development and deployment of endogenous AI, enabling large-scale intelligent distributed collaborative services. It compensates for the shortcomings of external computing power, promptly responding to changes in mobility and networks while fostering the emergence and development of future AI applications, such as immersive cloud XR, holographic communication, sensory interconnection, communication-sensing integration, and DT.
@@ 3.1.4 Security for AI
Some security vulnerabilities have remained after the design of 2G-5G systems, and conventional security measures have often failed to ensure safety. 6G networks necessitate establishing an inherently trustworthy and self-driving security system at the core of the system. Rapid advancements in cloud computing, big data, and AI technologies offer technical support to build a 6G network security system. In recent years, trusted endogenous security has emerged as a new security approach for 6G, characterized by four main features: collaboration, intelligent proactive defense, trustworthiness, and privacy protection. The concept of trusted endogenous security also applies to ensuring the security of AI.
(1) Ubiquitous Collaboration. Collaboration has become a technological approach to enhancing communication capabilities. Intelligent ubiquitous collaboration will enable NET4AI to adapt to various complex and dynamic environments and scenarios while improving the robustness of the AI plane to identify potential malicious attacks or abnormal behaviors accurately. Future security defenses will shift from isolated to highly efficient collaboration, achieving individual collaborative defenses between entities or intelligent agents, inter-layer defenses between protocol layers or architecture layers, and inter-domain defenses between network space domains.
(2) Intelligent Active Defense. The rapid development of 6G networks and AI technology presents opportunities for intelligent security defenses. The intelligent, proactive defense will enable NET4AI to continuously learn and adapt while automatically adjusting defense strategies based on changes in the network environment and AI security threats to counteract attacks promptly. This transition represents a shift from passive protection to proactive sensing and defense [143].
(3) Trustworthiness. 6G will further evolve towards trustworthiness capabilities, achieving a comprehensive endogenous security system encompassing security (i.e., confidentiality, integrity, availability), reliability, resilience, and safety [144]. This can provide a trusted ecosystem for NET4AI. Furthermore, flexible and reliable access control [145] and effective identity authentication[146] can, to some extent, prevent unauthorized access from impacting the endogenous trusted environment of NET4AI.
(4) Privacy Protection. The deep integration of AI, big data, and 6G networks will exacerbate the challenges of data privacy protection. NET4AI must handle massive business and personal data while preventing privacy leakage. Emerging distributed ML, while achieving distributed ubiquitous intelligence and partially avoiding privacy leakage caused by AI model training, still has security risks [147]. Therefore, NET4AI will integrate distributed ML with existing privacy protection methods at multiple points to enhance data privacy protection and build an efficient and secure data ecosystem.
@@ 3.1.5 Orchestration for AI
6G is considered to possess an extremely high level of network autonomy, which means that the management & orchestration capability of the network can conduct unified and dynamic orchestration and scheduling for various functional components of service requirements, thus enabling the self-optimization, self-evolution, and self-healing of the network. The 6G system will introduce external applications and service demands while exposing internal diversified resources and functions. This requires flexible cross-functional flow and scheduling of multi-dimensional resources, such as computing, communication, and storage within the network, offering customized and personalized application services based on virtualization and microservice technologies. For instance, AI capabilities and analyzed data within the network can be opened to third parties to provide services and necessary support, a crowdsourcing behavior that necessitates the support of management and orchestration functions. Furthermore, through joint analysis and prediction of computing resources and network traffic, containerized gNBs can automatically scale up or down based on predicted network traffic loads, enabling dynamic resource allocation and network energy savings. Therefore, management and orchestration capabilities are closely related to all other capabilities, orchestrating and managing resources and functions efficiently and flexibly across all functional layers of the 6G network, achieving joint optimization.
@@
Orchestration for AI of the intelligence plane signifies efficient support for deploying, operating, and optimizing AI services within the network through orchestration. It automatically configures and orchestrates network resources based on the specific requirements of AI services, ensuring that AI applications obtain the necessary computational power, storage, network bandwidth, and other resources. Additionally, this orchestration capability enables real-time monitoring of AI service performance, dynamically adjusting resource allocation to accommodate fluctuations in service loads, thereby guaranteeing the stability and efficiency of AI services [148]. This seamless integration and support facilitate smoother AI service operation within the network, providing robust network guarantees for various intelligent applications. Specifically, this manifests in the following three aspects:
@@
(1) Automated Deployment and Resource Allocation. The orchestration capability swiftly responds to AI services’ network resource demands through automation. Upon AI service deployment, the orchestrator intelligently analyzes the current network resource utilization, automatically configuring network devices, establishing links, and deploying services. This automated process enhances deployment efficiency, mitigates human errors, and ensures rapid AI service launch and stable operation.
@@
(2) Dynamic Optimization and Elastic Scaling. As AI services operate, their resource requirements may fluctuate. The orchestration capability features dynamic optimization and continuous monitoring of AI service performance, including CPU usage, memory consumption, network bandwidth, and other key metrics. Upon detecting resource shortages or surpluses, the orchestration system automatically adjusts resource allocation, such as adding compute nodes, expanding storage capacity, or optimizing network bandwidth, to meet AI services’ dynamic demands. Furthermore, the orchestration system supports elastic scaling, automatically resizing resources based on AI service loads, ensuring service stability and efficiency.
@@
(3) Fault Self-Healing and Intelligent O&M. By integrating monitoring, alerting, and troubleshooting functions, the orchestrator promptly identifies and locates faults within the network, automatically triggering recovery mechanisms. This fault self-healing capability minimizes service disruptions for AI services, enhancing service availability and reliability. Additionally, the orchestration facilitates intelligent network O&M, empowering network administrators to manage network resources and AI services more efficiently, reducing O&M costs, and improving efficiency.